OpenStudy (anonymous):

Find the derivative of f(x) = -6/x at x = 12.

OpenStudy (anonymous):

a/12?

sam (.sam.):

First set it \[f(x)=\frac{6}{x} \\ f(x)=6x^{-1}\] Then use power rule, power rule is \[\frac{d}{dx}(ax^n)=nax^{-1}\]

OpenStudy (anonymous):

derivative of f(x) = -6/x =-6x^-1 =-6(-1)x^-2 =6x^-2 =6/x^2 =6/144

sam (.sam.):

Don't forget the minus sign, \[f(x)=-\frac{6}{x}\]

OpenStudy (anonymous):

d/dx(6x^-1) = n(6)(1/x) = (6/x)n

OpenStudy (anonymous):

right -6/xn

OpenStudy (anonymous):

kitty where do i go from there?

OpenStudy (anonymous):

i simplify?

OpenStudy (anonymous):

to 1/24

OpenStudy (anonymous):

see remeber this formula wheverver u are having in this for \[x ^{n}\] then write in this form \[nxn ^{-1}\]

sam (.sam.):

That's the answer, but you will need to simplify that

OpenStudy (anonymous):

for=form

OpenStudy (anonymous):

do u still having problem ? :)

OpenStudy (anonymous):

okay I see. I have another one that is for derivative but it has the value of (t) in it. can u help?

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

lol ater answering urs i will do mine work that is study :)

OpenStudy (anonymous):

The position of an object at time t is given by s(t) = 1 - 10t. Find the instantaneous velocity at t = 10 by finding the derivative.

OpenStudy (anonymous):

sorry kid something happened to my computer now let me write :)

OpenStudy (anonymous):

oh lol no problem

OpenStudy (anonymous):

ds/dt=v s(t) = 1 - 10t ds/dt= d(1-10t)/dt ...........(differentiate it with respect to t) v=-10 ?????

OpenStudy (anonymous):

is that the answer ?

OpenStudy (anonymous):

1-10t d(1-10t)/dt = 1-10 = -9?

OpenStudy (anonymous):

see the derivative of constant is always zero here 1 is constant

OpenStudy (anonymous):

I dont know...

OpenStudy (anonymous):

write it now in this form d(1-10t)/dt = d(1)/dt - d(10t)/dt =0-10=-10 :)

OpenStudy (anonymous):

d(1-10t)/dt = d(1)/dt - d(10t)/dt = 0-10 = -10 :)

OpenStudy (anonymous):

kid see this example 2t^2 - 3t +3 differentiation of 2t^2 =2d(t^2)/dt = 2*2t=4t -d(3t)/dt=-3 d(3)/dt =0 :)simple :) u need practice :)

OpenStudy (anonymous):

So -10 is the answer?

OpenStudy (anonymous):

yeah i think so :) lol well i thi yes but still having doubt coz it's a simple question :O

OpenStudy (anonymous):

hey kid do u need software for help

OpenStudy (anonymous):

i studied that last yr (previous one)

OpenStudy (anonymous):

You doubt the answer because its so easy lol?

OpenStudy (anonymous):

yeah lol

OpenStudy (anonymous):

uhh okay you are on one lol. Thank you so much ! you deserve best answer x10

OpenStudy (anonymous):

kid i can give u list of site which will help u :) do u want ? lol hey kid today is mah first day :P

OpenStudy (anonymous):

i answer mostly on yahoo answers :) here is me lol http://in.answers.yahoo.com/activity?show=A7pnZeoWaa

OpenStudy (anonymous):

Ohhhh okay cool! I used to be on there!

OpenStudy (anonymous):

sites- http://easycalculation.com/

OpenStudy (anonymous):

oh u r familiar with taht site :)

OpenStudy (anonymous):

http://calculator.tutorvista.com/

OpenStudy (anonymous):

see there is solver for derivative also :)

OpenStudy (anonymous):

u can use wormalpha ,algebra solver just type ______ solver on search bar _____= derivative,limits, algebra ,can be anything

OpenStudy (anonymous):

@kittycat01 Thanks I see it now :D